The 3-month t bill is quoted at 1.5125-1.500 on a discount basis and has 86 days until maturity. How much would you pay for $10,000 face amount?
Answer #1

T-bills are quoted at a discount from face value, with the discount expressed as an annual rate based on a 360-day year.

Here the Face value F is $10,000 which means that after expiry i.e after 86 days the buyer of the bill would get $10,000 .

The question what is the Price P the buyer must pay so that the discount basis or annual yield would be 1.5125( this is the bid rate which the buyer is willing to offer) . By employing Formula of Simple interest F= P( 1+rT) This Implies P= F/( 1+rT)

P= 10000/[1+ 1.5125%*(86/360)] = $9963.99 0r $9964 (Ans)